Calculate the mean and the variance of the discrete random variable x which one values 12 and 3, given that P(1)=10/33, P(2)=1/3 and P(3)=12/33

1
Expert's answer
2022-03-31T07:55:59-0400

The mean:

"\mu=\sum x_i\cdot P(x_i)=\\
=1\cdot\cfrac{10}{33}+2\cdot\cfrac{1}{3}+3\cdot\cfrac{12}{33}=\cfrac{68}{33}\approx2.06."


The variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 1-\cfrac{68}{33}, 2-\cfrac{68}{33}, 3-\cfrac{68}{33}
\end{Bmatrix}="

"=\begin{Bmatrix}
\cfrac{-35}{33},\cfrac{-2}{33}, \cfrac{31}{33}
\end{Bmatrix},"

"\sigma^2=\\
=\begin{pmatrix} \cfrac{-35}{33}\end{pmatrix}^2\cdot\cfrac{10}{33}+\begin{pmatrix} \cfrac{-2}{33}\end{pmatrix}^2\cdot\cfrac{1}{3}+\begin{pmatrix} \cfrac{31}{33}\end{pmatrix}^2\cdot\cfrac{12}{33}=\\
=2.01."